please help need this one problem solved asap.
solve for x if the perimeter is 32.

one side of the weird looking triangle is 3,the other side is
x+7,and on the bottom is 2x-2 please someone help me..

Answer by mananth(16946)

one side of the weird looking triangle is 3,the other side is
x+7,and on the bottom is 2x-2 please someone help me..
side 1 3 units
side 2 x+7
Side 3 2x-2
3+x+x+7+2x-2 = 32
4x+8=32
4x=32-8=24
x=6